# Convergence and Generalization of Anti-Regularization for Parametric Models

### v1

@@ 4.2.2 CIFAR-10
Statistical comparisons in Table 11 show that the strength of ECE reduction varied across data splits. At the 0.5% split, the reduction was statistically significant ($p_{t}=1\times 10^{-4}$, Holm–Bonferroni adjusted $p_{t}^{\mathrm{Holm}}=0.0193$, $p_{W}=1\times 10^{-4}$). For the 0.3% split, the Wilcoxon test showed significance ($p_{W}=0.0099$), but the $t$-test failed to pass the adjusted threshold. For accuracy, no split yielded statistically significant differences, aligning with the observed near-zero changes. At the 0.2% split, since the baseline was optimal, no comparison was necessary.
Stability and safety diagnostics in Table 12 show that the projection intervention rate $r_{\mathrm{proj}}$ was 100.0 across all splits, indicating the trust-region constraint remained consistently active. The gradient clipping rate $r_{\mathrm{clip}}$ was 0% for SGDM (used at 0.1–0.2%) but high (89–94%) for Adam (used at 0.3–1.0%), suggesting optimizer-specific variations in the activation of safety guards while maintaining stable training. The output scale diagnostic $\rho$ ($\times 10^{-3}$) peaked at 98.7 for the 0.1% split but remained within the safe range; other splits showed lower values between 3.3 and 45.3. Epochs and runtime scaled gradually with increasing data size.
@@ 4.3 Ablation Study
Figure 1 presents the performance curves (with 95% confidence intervals) across different $\alpha$ values for each dataset. In regression tasks (UCI Concrete and Airfoil), the RMSE curves remain nearly flat over the range $\alpha\in[0.5,1.5]$, and differences between values fall within the confidence intervals. This indicates low sensitivity to $\alpha$, consistent with our theoretical recommendation of $\alpha\approx 1$ for regression. In contrast, classification tasks (MNIST and CIFAR-10) exhibit mild but directional trends. On MNIST, accuracy slightly decreases with increasing $\alpha$ when using Adam, while SGDM shows a minor advantage at lower $\alpha$ values. On CIFAR-10, Adam achieves near-optimal performance around $\alpha\approx 0.5$, whereas SGDM shows gradual improvement with larger $\alpha$. These results suggest that $\alpha\approx 0.5$ is a reasonable starting point for classification, and $\alpha\approx 1$ remains robust for regression.
Table 13 compares the effects of Safety ON vs OFF at the reference $\alpha$ values. The outcomes diverge by task difficulty and optimizer. For regression, SGDM suffers from numerical divergence when Safety is OFF, with RMSE spiking to $10^{8}\sim 10^{10}$, while Safety ON (trust-region + gradient clipping) stabilizes training. Adam avoids divergence but still incurs RMSE increases of about $0.6$ to $0.7$, indicating a modest cost from conservative safety guards. For classification, MNIST shows a minor performance drop under Safety ON (accuracy reduced by 0.0010.0010.0070.007), implying near-neutral or slightly adverse effects. However, on the more challenging CIFAR-10, both Adam and SGDM gain from Safety ON with accuracy improvements of +0.018+0.018 and +0.025+0.025 respectively, suggesting stronger benefits of safety in high-difficulty settings.
Table 14 reports results under the best-performing decay exponent $\alpha_{\text{best}}$, reaffirming these trends. In regression, SGDM again diverges under Safety OFF (e.g., Airfoil RMSE reaching $2.51\times 10^{14}$), confirming Safety as a practical necessity. Adam remains slightly inferior in RMSE. For classification, MNIST still shows only marginal losses (Adam: 0.010-0.010, SGDM: 0.0002-0.0002), while CIFAR-10 benefits consistently from Safety ON (Adam: +0.018+0.018, SGDM: +0.025+0.025).
(a) UCI Concrete (b) UCI Airfoil
Refer to caption Refer to caption
(c) MNIST (d) CIFAR-10
Refer to caption Refer to caption
Figure 1: Performance curves versus decay exponent $\alpha$. Shaded regions denote 95 % confidence intervals over eight seeds.
@@ 5 Discussion
Across all experiments, Anti-regularization (AR) mitigated underfitting in low-data regimes, improving performance, and gradually faded as data increased, converging to the baseline ($\lambda=0$) under the power-law decay schedule. The mechanism acted helpfully where expressive power was needed, and neutrally where it was not.
Mechanism: Separation of accuracy and calibration.
In classification, the role of AR varied by task difficulty. On MNIST, where both accuracy and calibration were already high, AR had little effect. On CIFAR-10, AR consistently reduced ECE while preserving accuracy, suggesting its role in improving calibration without harming accuracy.
Effects and stability in regression.
@@
On both UCI Concrete and Airfoil, AR consistently reduced RMSE and improved $R^{2}$ in small to medium data regimes. SGDM was highly unstable without safety components, but trust-region constraints, gradient clipping, and projection restored stable convergence. Adam was more inherently stable, though safety guards incurred minor performance costs, which diminished as $\lambda$ decayed with more data.
Sensitivity to the decay exponent $\alpha$.
@@
Ablation results aligned with theory: regression was robust and flat around $\alpha\approx 1$, while classification favored $\alpha\approx 0.5$. Thus, fast decay suits regression, and slow decay suits classification.
Interpretation of safety conditions and diagnostics.
@@
The implemented safety mechanisms ensured overall stability. Projection rates were high, indicating consistent adherence to trust regions, while clipping rates adapted to optimizer and data regime. The output scale ratio $\rho$ mostly stayed near 1, and rare large values did not indicate divergence when interpreted in absolute scale. Safety guards were essential under risky settings and imposed minimal costs otherwise.
Optimizer dependency and practical implications.
In regression, SGDM was highly sensitive to safety settings and should default to safety ON. Adam was more forgiving, though tuning trust-region radius and clipping threshold conservatively helped minimize cost. In classification, both optimizers maintained accuracy, and calibration gains were more pronounced in harder datasets.
Statistical strength and reproducibility.
@@
Paired comparisons across multiple seeds confirmed consistent effect directions, and confidence intervals narrowed rapidly with more data. Some improvements did not survive conservative corrections, but were reproducible across multiple splits. All logs ($\lambda$, $\rho$, $r_{\mathrm{clip}}$, $r_{\mathrm{proj}}$) and configurations are publicly released to support independent verification.
Limitations and future work.
The study focused on compact models and controlled settings; thus, the optimal $\alpha$ and safety margins may vary for deeper architectures and higher-dimensional spaces. Future work includes evaluating AR in distribution shift settings, exploring alternatives such as degree-of-freedom targeting, and systematically analyzing interactions with learning rate, batch size, weight decay, and early stopping.
Practical guidance.
For regression, default to $\alpha\approx 1$ and always enable safety with SGDM. With Adam, activate safety selectively and minimize cost by narrowing trust regions and clip thresholds. For classification, explore $\alpha\in[0.25,0.75]$, and default to safety ON for harder tasks to improve calibration while preserving accuracy. In general, projection and clipping act as low-cost safeguards and can remain always active.
@@ 6 Conclusion
This study introduced Anti-regularization, a loss-based intervention that alleviates underfitting in small-data regimes and naturally fades with increasing sample size via a decay schedule. We established spectrum- and trust-region-based safety conditions, and designed a safeguard that combines projection and gradient clipping to block divergence and nullify intervention when unnecessary. Experiments showed that faster decay is suitable for regression and slower decay for classification, with persistent safety activation benefiting harder tasks. In effective regimes, the method boosted expressivity and generalization; in less effective ones, it reverted smoothly to the base model without degrading performance. From an implementation perspective, the simple decay schedule and lightweight safety components were easily integrated into standard pipelines. Future work includes optimizing the decay exponent for larger models and expressive architectures, validating robustness and calibration under distribution shifts, quantitatively comparing alternative schedules like degree-of-freedom targeting, and analyzing interactions with learning rate, batch size, weight decay, and early stopping. Our study provides reproducible code and procedures, offering a practical guide for reliable learning under constrained data and resource conditions.
@@ A Fast CI run (one seed/fraction):
Appendix A Preliminaries and Notation
A.1 Notation Summary
Throughout this appendix, we assume a fixed design where the sample size is denoted by $n:=|S|$. Unless otherwise specified, expectations $\mathbb{E}[\cdot]$ are taken with respect to the noise variable $\varepsilon$, conditional on the inputs.
@@ A.1 Notation Summary
Symbol Meaning Shape/Domain $S=\{(x_{i},y_{i})\}_{i=1}^{n}$ Training dataset $x_{i}\in\mathbb{R}^{p},\ y_{i}\in\mathbb{R}\ \text{or}\ \{1,\dots,C\}$ $X=[x_{1}^{\top};\dots;x_{n}^{\top}]$ Design matrix $\mathbb{R}^{n\times p}$ $y=(y_{1},\dots,y_{n})^{\top}$ Response vector $\mathbb{R}^{n}$ $\theta$ Parameter vector (or network weights) $\mathbb{R}^{p}$ or multi-tensor $f_{\theta}(\cdot)$ Prediction function $\mathbb{R}^{p}\to\mathbb{R}$ or $\to\Delta^{C-1}$ $\ell(\theta;x,y)$ Loss function Reg: squared loss; Cls: cross-entropy $\hat{F}_{\lambda}(\theta)$ Regularized empirical risk objective $\frac{1}{n}\sum\ell(\theta;x_{i},y_{i})-\lambda R(\theta)$ $\lambda$ Anti-regularization coefficient (reward strength) $\lambda\geq 0$ $R(\theta)$ Regularization/reward term $\tfrac{1}{2}\|\theta\|_{2}^{2}$ or $\tfrac{1}{2}\theta^{\top}W\theta$ $W\succeq 0$ Spectrum-weighting matrix $\mathbb{R}^{p\times p}$ $\hat{\Sigma}=\tfrac{1}{n}X^{\top}X$ Sample covariance (fixed design) $\mathbb{R}^{p\times p}$ $\sigma_{\min/\max}(A)$ Minimum/maximum eigenvalue of symmetric AA $\mathbb{R}_{\geq 0}$ $S_{\lambda}$ Linear smoother $X\big{(}X^{\top}X-n\lambda I\big{)}^{-1}X^{\top}\in\mathbb{R}^{n\times n}$ $\operatorname{df}(\lambda)$ Degrees of freedom $\operatorname{tr}(S_{\lambda})$ $\varepsilon$ Noise $\mathbb{E}[\varepsilon]=0,\ \operatorname{Var}(\varepsilon)=\tau^{2}I$ $\tau^{2}$ Noise variance (homoskedastic) $\mathbb{R}_{>0}$ $H(p_{y},q_{\theta}(x))$ Classification loss (CE) $p_{y}$: one-hot, $q_{\theta}$: predicted distribution $m_{\theta}(x,y)$ Margin Defined per problem via logit differences, etc. $\phi$ Bounded Lipschitz reward function Bounded and Lipschitz $G$ Network Lipschitz constant $\mathbb{R}_{>0}$ $K$ NTK/kernel matrix $\mathbb{R}^{n\times n}$ $\|\cdot\|_{\mathcal{H}_{K}}$ RKHS norm Induced by kernel $K$
Assumptions of the Regression Model
We assume the model $y=X\theta^{\star}+\varepsilon$, where $\varepsilon\sim(0,\tau^{2}I)$. The estimator is expressed via the linear smoother as $\hat{f}_{\lambda}=S_{\lambda}y$.
Definition of Empirical and Test Risk
We consistently adopt the following definitions throughout the paper:
@@
where $y^{\text{new}}=X\theta^{\star}+\varepsilon^{\prime}$, $\varepsilon^{\prime}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}(0,\tau^{2}I)$, and $\varepsilon^{\prime}\perp\!\!\!\perp\varepsilon$. Under these definitions, the optimism identity holds as
@@
For a detailed derivation, see Section A.3.
A.2 Basic Spectral Facts
This subsection collects standard facts repeatedly used in the subsequent proofs. Unless stated otherwise, all matrices are assumed to be real and symmetric.

### v2

@@ 5.2.2 CIFAR-10
Statistical comparisons (Table 11) show that the strength of calibration improvements varied across splits. At the 0.5% split, the reduction in ECE was statistically significant ($p_{t}=1\times 10^{-4}$, Holm–Bonferroni adjusted $p_{t}^{\mathrm{Holm}}=0.0193$, $p_{W}=1\times 10^{-4}$). At the 0.3% split, ECE improvements were significant only under the nonparametric test ($p_{W}=0.0099$), while the corrected $t$-test did not reach significance. For accuracy, no significant differences were detected at any split, consistent with the observation that accuracy changes were practically negligible. At the 0.2% split, the Baseline was optimal and thus no comparison was applicable.
Stability and safety diagnostics (Table 12) indicate that projection operator intervention rate $r_{\mathrm{proj}}$ remained at 100.0 across all splits, confirming that trust-region constraints were always active. Gradient clipping rate $r_{\mathrm{clip}}$ was 0% for SGDM at the 0.1–0.2% splits, but much higher (89–94%) for Adam at the 0.3–1.0% splits. This suggests that the pattern of safeguard activation varied with the optimizer and sample fraction, yet training remained stable throughout. The output-scale safety ratio $\rho$ (scaled by $10^{-3}$) was highest at 0.1% (98.7) but still within the safe region, and ranged between 3.3 and 45.3 for other splits. Epoch counts and runtime scaled moderately upward with larger sample fractions, consistent with expected patterns.
@@ 5.3 Ablation Study
Tables 13 and 14 summarize performance across values of the decay exponent $\alpha$ as mean $\pm$ 95% confidence intervals. In regression (UCI Concrete and Airfoil), RMSE differences across $\alpha\in\{0.5,1.0,1.5\}$ were minimal, with confidence intervals largely overlapping, indicating low sensitivity to $\alpha$. This aligns with our theoretical guidance that $\alpha\approx 1$ is appropriate for regression tasks. In classification (MNIST and CIFAR-10), MNIST with Adam showed virtually no dependence on $\alpha$, while MNIST with SGDM exhibited a slight downward trend as $\alpha$ increased, though the changes were within confidence bounds. For CIFAR-10, Adam achieved its best performance around $\alpha\approx 0.5$, while SGDM peaked modestly at $\alpha=0.75$, but no consistent monotonic trend was observed. In short, classification tasks suggest $\alpha\approx 0.5$ as a reasonable starting point, while regression remains robust near $\alpha\approx 1$.
Table 15, which compares the effect of Safety at the reference decay exponent, shows contrasting patterns depending on task difficulty and optimizer. In regression, SGDM diverged numerically with Safety OFF, producing RMSE on the order of $10^{8}\sim 10^{10}$, but returned to stable convergence with Safety ON (trust-region constraint + gradient clipping). Adam did not diverge, but incurred a modest RMSE increase of about 0.60.70.6-0.7, reflecting a small cost of conservative safeguards. In classification, MNIST exhibited slight disadvantages under Safety ON, with accuracy decreases of 0.0010.0070.001-0.007, effectively neutral to mildly adverse. By contrast, CIFAR-10 benefited from Safety ON, with accuracy improvements of +0.018+0.018 and +0.025+0.025 for Adam and SGDM, respectively, suggesting that the net benefit of Safety grows with task difficulty.
The same pattern persists at the best decay exponent $\alpha_{\text{best}}$ (Table 16). For regression, SGDM again diverged severely without Safety (e.g., Airfoil RMSE of $2.51\times 10^{14}$), reaffirming that Safety is essentially a necessary condition, while Adam remained slightly disadvantaged in RMSE under Safety ON. For classification, MNIST still showed only marginal losses (Adam 0.010-0.010, SGDM 0.0002-0.0002), whereas CIFAR-10 consistently gained from Safety ON with both optimizers (Adam +0.018+0.018, SGDM +0.025+0.025).
Table 13: Regression performance at decay exponents $\alpha\in\{0.5,\,1.0,\,1.5\}$. Entries are mean $\pm$ 95% confidence intervals over seeds.
Dataset Optimizer $0.5$ $1.0$ $1.5$
[…]
SGDM $39.497\,\pm\,1.772$ 39.389± $\pm\,1.744 39.398$± $398\,\pm\,1.67$8
UCI Airfoil Adam $124.642\,\pm\,1.141$ $124.573\,\pm\,1.128$ $124.634\,\pm\,1.098$
[…]
Table 14: Classification performance at decay exponents $\alpha\in\{0.25,\,0.5,\,0.75\}$. Entries are mean $\pm$ 95% confidence intervals over seeds.
Dataset Optimizer $0.25$ $0.5$ $0.75$
[…]
SGDM $0.9391\,\pm\,0.0041$ $0.9368\,\pm\,0.0049$ $0.9351\,\pm\,0.0046$
CIFAR-10 ADAM 0.4410± $410\,\pm\,0.011$1 $0.4460\,\pm\,0.0122$ $0.4452\,\pm\,0.0138$
SGDM $447\,\pm\,0.0134 0.440$$402\,\pm\,0.0086$ $0.4473\,\pm\,0.0124$
@@ 6 Discussion
Across all experiments, Anti-regularization mitigated underfitting in small-sample regimes, improving performance, while its intervention effects naturally vanished under the power-law decay schedule as sample size increased, converging to the Baseline ($\lambda=0$). In other words, it acted as a procedure that provided assistance when additional expressivity was needed, but became harmless in stable regimes.
Mechanism: Separation of Accuracy and Calibration
In classification, the contribution of AR varied with task difficulty. On MNIST, where both accuracy and calibration were already high, AR interventions were effectively neutralized. On CIFAR-10, accuracy was maintained while ECE consistently decreased. This suggests that in classification AR is particularly sensitive to suppressing overconfidence, improving calibration while remaining neutral with respect to accuracy.
Effects and Stability in Regression
@@
For both UCI Concrete and Airfoil, repeated improvements in RMSE reduction and $R^{2}$ increase were observed in small- to mid-sized sample regimes. With SGDM, training was prone to divergence without safeguards, but enabling trust-region constraints, gradient clipping, and projection operators immediately restored stable convergence. Adam was inherently more stable, though conservative safeguards sometimes incurred slight performance costs, which diminished quickly as $\lambda$ decayed with increasing sample size.
Sensitivity to the Decay Exponent $\alpha$
@@
Results from the ablation study aligned with theoretical recommendations. Regression was flat and robust around $\alpha\approx 1$, while classification showed $\alpha\approx 0.5$ as a reasonable starting point. Thus, faster decay is appropriate for regression, while slower decay is suitable for classification, confirming the proposed guidance.
Interpretation of Safety Conditions and Diagnostics
@@
Safety conditions implemented via projection and clipping ensured overall convergence stability. Projection intervention rates were relatively high, indicating that updates consistently respected trust-region constraints, while clipping rates adapted according to optimizer and sample regime. The output-scale ratio $\rho$ remained mostly near unity, and even the rare large values did not indicate runaway behavior when absolute scales were considered. Taken together, safeguards were essential in high-risk configurations, and in low-risk ones they imposed only minor costs while remaining effectively suppressed.
Optimizer Dependence and Practical Implications
In regression, SGDM was highly sensitive to the presence of safeguards, making it advisable to enable them by default. Adam was comparatively tolerant, but conservative choices of trust-region radius and clipping thresholds reduced its performance costs. In classification, both optimizers generally maintained accuracy, with calibration benefits being more pronounced on more challenging datasets.
Statistical Strength and Reproducibility
@@
Consistency in the direction of effects was confirmed through seed-paired comparisons across conditions, with confidence intervals narrowing rapidly as sample size increased. Some moderate improvements did not survive conservative multiple-comparison corrections, but the underlying phenomena were reproduced across multiple splits. All logs ($\lambda$, $\rho$, $r_{\mathrm{clip}}$, $r_{\mathrm{proj}}$) and configurations are released in the repository, enabling independent verification.
Limitations and Future Work
By focusing on simple models under controlled settings, our study leaves open the possibility that optimal $\alpha$ values and safe radii differ for more complex architectures and high-dimensional representation spaces. Future work should extend quantitative validation to distribution shift scenarios in calibration and robustness, explore alternative decay rules such as degrees-of-freedom targeting, and systematically examine interactions with learning rate, batch size, weight decay, and early stopping.
Practical Guidelines
For regression, $\alpha\approx 1$ is recommended as a default, with safeguards always enabled when using SGDM. For Adam, safeguards may be applied selectively depending on task and split, but should use conservative trust-region radii and clipping thresholds to minimize costs. For classification, a light search within $\alpha\in[0.25,\,0.75]$ is appropriate. On more difficult datasets, safeguards should be activated by default to achieve calibration improvements while maintaining accuracy. Overall, projection and clipping act as conservative guardrails that prevent failure at little cost and can thus be safely kept enabled at all times.
@@ 7 Conclusion
This study introduced Anti-regularization, which augments the loss function with a reward term to mitigate underfitting in small-sample regimes, while employing a decay schedule that ensures the intervention naturally weakens as sample size increases. We established safety conditions based on spectral and trust-region criteria, and combined projection operators with gradient clipping to prevent divergence, with interventions automatically vanishing when ineffective. Across experiments, regression favored relatively fast decay, while classification benefited from slower decay, with continuous activation of safeguards contributing to stable learning in more challenging tasks. Where effective, Anti-regularization amplified expressivity to aid generalization; where negligible, it naturally regressed to the baseline model without harming performance. From an implementation perspective, we confirmed that Anti-regularization integrates seamlessly into existing pipelines with only simple schedules and lightweight safeguards. Future directions include exploring optimal decay exponents in large-scale models and complex representation spaces, validating robustness and calibration under distribution shifts, quantitatively comparing alternative schedules such as degrees-of-freedom targeting, and analyzing interactions with hyperparameters such as learning rate, batch size, weight decay, and early stopping. By releasing reproducible code and procedures, this work supports adoption and extension in practice, providing practical guidance for reliable learning under limited data and resource constraints.
@@ A Fast CI run (one seed/fraction):
Appendix A Preliminaries and Extended Notation
A.1 Summary of Notation
Throughout this appendix, the sample size $n:=|S|$ is considered under a fixed design setting. Unless otherwise specified, expectations $\mathbb{E}[\cdot]$ are taken with respect to the noise $\varepsilon$.
@@ A.1 Summary of Notation
Symbol Meaning Shape/Range $S=\{(x_{i},y_{i})\}_{i=1}^{n}$ Training dataset $x_{i}\in\mathbb{R}^{p},\ y_{i}\in\mathbb{R}\ \text{or}\ \{1,\dots,C\}$ $X=[x_{1}^{\top};\dots;x_{n}^{\top}]$ Design matrix $\mathbb{R}^{n\times p}$ $y=(y_{1},\dots,y_{n})^{\top}$ Response vector $\mathbb{R}^{n}$ $\theta$ Parameter vector (or network weights) $\mathbb{R}^{p}$ or multi-tensor $f_{\theta}(\cdot)$ Predictor function $\mathbb{R}^{p}\to\mathbb{R}$ or $\to\Delta^{C-1}$ $\ell(\theta;x,y)$ Loss function Regression: squared loss; Classification: cross-entropy, etc. $\hat{F}_{\lambda}(\theta)$ Empirical risk with reward term $\tfrac{1}{n}\sum\ell(\theta;x_{i},y_{i})-\lambda R(\theta)$ $\lambda$ Anti-regularization coefficient (reward strength) $\lambda\geq 0$ $R(\theta)$ Regularization/reward term $\tfrac{1}{2}\|\theta\|_{2}^{2}$ or $\tfrac{1}{2}\theta^{\top}W\theta$ $W\succeq 0$ Spectral weighting matrix $\mathbb{R}^{p\times p}$ $\hat{\Sigma}=\tfrac{1}{n}X^{\top}X$ Sample covariance (fixed design) $\mathbb{R}^{p\times p}$ $\sigma_{\min/\max}(A)$ Minimum/maximum eigenvalue of symmetric AA $\mathbb{R}_{\geq 0}$ $S_{\lambda}$ Linear smoother $X\big{(}X^{\top}X-n\lambda I\big{)}^{-1}X^{\top}\in\mathbb{R}^{n\times n}$ $\operatorname{df}(\lambda)$ Degrees of freedom $\operatorname{tr}(S_{\lambda})$ $\varepsilon$ Noise $\mathbb{E}[\varepsilon]=0,\ \operatorname{Var}(\varepsilon)=\tau^{2}I$ $\tau^{2}$ Noise variance (homoskedasticity) $\mathbb{R}_{>0}$ $H(p_{y},q_{\theta}(x))$ Classification loss (CE) $p_{y}$: one-hot, $q_{\theta}$: predictive distribution $m_{\theta}(x,y)$ Margin Defined per task (e.g., logit differences) $\phi$ Bounded margin reward function Bounded and Lipschitz $G$ Network Lipschitz constant $\mathbb{R}_{>0}$ $K$ NTK/kernel matrix $\mathbb{R}^{n\times n}$ $\|\cdot\|_{\mathcal{H}_{K}}$ RKHS norm Corresponding to kernel $K$
Basic Assumptions of the Regression Model
We assume $y=X\theta^{\star}+\varepsilon,\ \varepsilon\sim(0,\tau^{2}I)$. The estimator is expressed as a linear smoother $\hat{f}_{\lambda}=S_{\lambda}y$.
Definitions of Empirical and Test Risk
Throughout this paper, we consistently use the following definitions:
@@
where $y^{\text{new}}=X\theta^{\star}+\varepsilon^{\prime}$, $\varepsilon^{\prime}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}(0,\tau^{2}I)$, and $\varepsilon^{\prime}\perp\!\!\!\perp\varepsilon$. Under this definition, the optimism identity holds as
@@
For detailed derivation, see §A.3.
A.2 Basic Facts on Spectral Matrices
We collect here standard facts that are repeatedly used in subsequent proofs. Unless otherwise specified, all matrices are real symmetric.